7) Find the derivative of the function at P. (-7,-8) in the direction of ū= 12i – 5j for f(x,y) = tan-1 -4633)
Answer #1

fcx,y) = tan (3) 32 f(x,y) tañ ( = a οχ (4) It (139) y² 6) 179202 - 34 1+902 D fyxz, y) = oy tao( 9 ) (**) y² 179002 3x 2 14- 129-55 given Cool vert to Unit vectoon (121-5) C Q u, [44 +25 129 57 13 - le 13 여름 a = 72. Hene 13 pilafza +ly to - ) - (유다